# Kang M, Chen D, Pu Z, Gao J, Yu W. Spatio-Temporal Graphical Counterfactuals: An Overview

### v2

@@ 4 Differences between POM and SCM
In this section, the fundamental differences between POM and SCM are discussed, including (i) counterfactual falsifiability: POM assumes the positivity, thus the counterfactuals in POM can be falsified on observations, but SCM does not. The counterfactuals in SCM can be outside the observations, thus it is unfalsifiable. (ii) using back-door criterion to satisfy ignorability: POM requires ignorability to guarantee the unconfoundedness and correctness of counterfactual inference, but it is a complex human-guided operation. By contrast, SCM realizes unconfoundedness by the causal sufficiency and back-door adjustment, that are equivalent to the ignorability assumption. SCM facilitates counterfactual inference through graphical language. (iii) direct causation and indirect causation: POM struggles to process the indirect causation, and it typically relies on Baron-Kenny model. But in contrast, SCM has advantages to adjust the front-door path (that has indirect causation) through equivalently applying the back-door criterion twice. We further use this advantage to design the forward counterfactual inference algorithm to infer the spatio-temporal counterfactuals autonomously. The following are the details.
@@ 4.1 Counterfactual Falsifiability
As shown in Fig. 3, if we query to a POM, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”, we must ensure there are at least two observational samples, and they have the same values $X_{1}=0.5$ but different values in $X_{2}$. That is, one is $X_{2}=1$, and another one is $X_{2}=2$, to satisfy the positivity assumption (see Assumption 2.2). And then, due to the consistency assumption (see Assumption 2.2), we can infer the counterfactuals through matching the two samples exactly. But this is different in SCM, we can infer the counterfactuals through the Pearl’s causal ladder, even if we only have a single sample, as shown in Section 3.1. Thus, how to falsify the counterfactuals?
@@
It is a difficult question, because the counterfactuals are unobserved according to the definition (see “?” in Table 1), and we need at least two samples for comparison. In general, we cannot back to the past to falsify it with a different treatment, and we cannot reproduce the same experiments completely at different times. Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [7].
@@
Another explanation is that, without the positivity, but with the consistency only in POM, the counterfactuals can be extrapolated by regressive approaches, and they are falsifiable. For example, as shown in Fig. 5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$. Then, we query “if had done $T=4$, what would Y=?Y=?”, the counterfactual answer is $Y=2.5$, and the answer can be falsified by the regressive function because it does not beyond the observation. Thus, the consistency assumption actually claims that the counterfactuals can be falsified by the observations. This is also true for SCM, because SCM also uses the regressive approaches to infer counterfactuals.
@@
Moreover, there may be some confusions regarding the preset Bayesian network (see Fig. 2). To build an SCM, a Bayesian network is necessary to be known first. However, how to discover a probabilistic network from a single sample? This is usually impossible. A reasonable explanation is that, the network structure can be discovered from the system mechanisms (e.g., physical mechanisms, communication connections), or graphical knowledge learned from other domains, instead of probability. However, if it is accepted, the definition of causality could be broader and more ambiguous. This is why probabilistic and causal Bayesian network are often indistinguishable, as stated in Pearl’s book [7].
@@
Thus, actually, SCM provides a way to think the counterfactuals, and it emphasizes the thinking way is counterfactual, but not the thinking result, because the result cannot always be falsified without the positivity and consistency assumptions. And on the contrary, POM emphasizes the thinking result is counterfactual. This is a fundamental difference in the problem of counterfactual falsifiability.
@@ 4.2 Ignorability and Back-door Criterion
In the view of statistics, the ignorability assumption (see Assumption 2.2) is equivalent in calculation to the causal sufficiency assumption (see Assumption 3.1) with a back-door criterion. The accurate definition of back-door criterion can be found in these papers and books [17, 18, 4, 5]. Here we present another from with respect to the Table 1, defined as
@@
In Table 1, the covariate $X$ is said to satisfy back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
@@
Figure 4: A diagram of data extrapolation.
Figure 5: A diagram of back-door path.
@@
which is the same in case of $do(T=0)$ [17, 18, 4, 5]. Then, if the causal sufficiency is satisfied, that is, no hidden variable can interfere $X,Y,T$, the expectation of potential outcomes in Table 1 can be calculated by
@@
This is the same in POM, as follow:
@@
which is the same in case of $do(T=0)$. Thus, with comparing Eq. (9) and Eq. (10), $\mathbb{E}[Y|do(T=1)]=\mathbb{E}[Y(1)]$, and $\mathbb{E}[Y|do(T=0)]=\mathbb{E}[Y(0)]$, in the two frameworks, but they start from different assumptions. Thus, the ignorability assumption can be decomposed as the causal sufficiency assumption and the back-door criterion, and the causal sufficiency is relatively trivial to satisfy. Actually, the ignorability is more likely to be a technical assumption, that needs many controlled trials to deconfound, while the back-door path is clear to search in a Bayesian network. A potential and more difficult technical issue is to discover the network structure accurately.
@@ 4.3 Direct and Indirect Causation
Suppose that there is a causal ordered pathway from one treatment variable to another outcome variable. Then, if there are one or more variables between the two endpoints, the pathway is called as indirect causation, otherwise, it is called as directed causation. Actually, the indirect causation is also called as mediating effect, that has been discussed deeply in Rubin’s papers [74, 75, 76]. Here, Baron-Kenny model [19, 20] is introduced to present mediating analysis for potential outcome with graph. As shown in Fig. 7, $T\rightarrow Y$ is direct causation, and $T\rightarrow X\rightarrow Y$ is indirect causation.
Figure 6: A diagram of mediating analysis [19]. Here, $a,b,c$ are coefficients and $U_{X},U_{Y}$ are biases. $b^{\prime}$ and $U_{total}$ are the coefficient and bias respectively for total effect.
Figure 7: A diagram of front-door adjustment [5]. Here, $T$ is a treatment variable, and $Y$ is a target variable. $X_{1}$ and $X_{2}$ are both covariates.
To analyse indirect causation, causal-steps approach [19] is proposed, through building three linear regressive models, as shown in Fig. 7. Suppose that the regressive models are fitted sufficiently, the indirect causation would be detected if satisfy
@@
Another approaches are to test the significance of $H_{0}:b^{\prime}-b=0$ [77, 78, 79]. Or, to test the significance of $H_{0}:ab=0$ [80, 81]. If the correct regressive models are built, the counterfactuals can be calculated, just like SCM in Fig. 2. Moreover, they are all regression-based approaches, that restrict the usage in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and the test process would be time-consuming seriously.
But on the contrary, the test process can be faster with Bayesian network in an SCM. Here, we propose an algorithm to realize forward counterfactual inference as shown in Algorithm 1. To clearly demonstrate the principle of Algorithm 1, we first introduce the front-door criterion as follows:
@@
The covariate $X$ is said to satisfy front-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
@@
Then, a classic example is provided to show the front-door adjustment through the theorem 3.4.1 in book [5], as also shown in Fig. 7 here. If given a treatment variable $T$ and a target variable $Y$, with two covariates $X_{1}$ and $X_{2}$. Then one can adjust the distribution $P(Y|do(T))$ with the front-door criterion as follows:
@@
    If there are one or multiple ordered causal pathways from $T$ to $Y$ and there is at least one mediator on the ways, then the causal pathway can be decomposed into two causal pathways through that mediator, e.g., $T\rightarrow X_{1}$ and $X_{1}\rightarrow Y$ in Eq. (11).
@@
Thus, the front-door principle is non-atomic and decomposable. And with this more foundational principle, we propose the forward counterfactual algorithm to autonomously apply this adjustment principle, as presented in Algorithm 1. The algorithm can be conducted on arbitrary variable set $\{X_{1},\dots,X_{n}\}$ that satisfied causal sufficiency. And we also provide an example to show the inference process as shown in Fig. 8.
In Algorithm 1, the approaches to realize the codes in lines 1 and 3 can be various, e.g., some variants of depth-first-search algorithm and breadth-first-search algorithm [82]. Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set. Meanwhile, if the computational cost of factorization and adjustment are overlooked, the time complexity of the code in line 9 in Algorithm 1 is $\mathcal{O}(|\mathbf{V}|)$, because we only need to traverse all mediators in each pathway forward, and this process can be performed in parallel. Thus, the total time complexity of Algorithm 1 is $\mathcal{O}(|\mathbf{V}|^{2}+|\mathbf{V}||\mathbf{E}|)$. Thus, if the network structure of the Bayesian network is large, and it can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
5 Spatio-Temporal Graphical Counterfactuals
Then for the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns? And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? And, would someone still purchase the product even if they had never been shown the advertisement? In these real-world scenarios, interactive behaviors with lagged causal effects are widely-existing. POM cannot answer these questions, because POM does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 2.2), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [83, 84] and Differences in Differences [85, 86]). And this is the same to SCM, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes. Thus, to answer these counterfactual queries for interactive units and lagged causal effects, a concept of spatio-temporal graphical counterfactuals is proposed here.
@@ 5 Spatio-Temporal Graphical Counterfactuals
8      For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to back-door criterion (see Definition 4.1),
@@
10
11
12Otherwise, for $q$ pathways starting from $X_{j_{p}}$, select $q$ first mediators $X_{j_{p_{1}}},\dots,X_{j_{p_{q}}}$. Then, a group of source-mediator pairs, e.g., $(X_{j_{p}},X_{j_{p_{q}}})$, can be obtained;
13Factorize $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ as
@@
14Solve each distribution $P(\cdot|do(\cdot),\dots,do(\cdot))$ in Eq. (12) by Algorithm 1 recursively.
@@
Figure 8: An example of Forward Counterfactual Inference algorithm. Here, $X_{i}$ is the same as $X_{i}^{\prime}$ to distinguish two recursions.
@@
5.1 Spatio-Temporal Bayesian Networks
Before the buildings of SCM, actually, various causal graphical models have been proposed to model the spatio-temporal causality. As shown in Fig. 9, Temporal Bayesian Network [59, 87] are proposed to model the momentary causal dependency in a first-order Markov process. Further, Dynamical Bayesian Networks [59] are proposed to represent a combination of multiple Temporal Bayesian Networks. Full Time Graph [6] is proposed to summarize the temporal Bayesian dependency at full timestamps. Note that, Full Time Graph allows the instantaneous causal effects between two units (or variables) in a high-order Markov process, but it only allows strictly stationary causality, that is, the causal dependencies do not change with time. Moreover, Full Time Graph is proved to uniquely decomposed into multiple high-order Temporal Bayesian Networks, named UCN, if the instantaneous causal effects is not allowed [21]. The comparisons of these causal graphs are shown in Table 2.
Figure 9: Diagrams of various causal graphs.
Thus, based on the characteristics of these causal graphs, STBNs are defined as a group of DAGs with different timestamps, as shown in Fig. 9. In STBNs, intuitively, the nodes are not equivalent to the variables, but represent the state of the variables at some time steps. With the assumptions of causal Markov (see Assumption 5) and faithfulness (see Assumption 6), the joint distribution of all nodes can be factorized as
$P(X_{1,t},\dots,X_{n,t},X_{1,t-1},\dots,X_{n,t-T+1})$ (13)
$=$ $\prod_{\tau=0}^{T-1}\prod_{i=1}^{n}P(X_{i,t-\tau}|\mathbf{Pa}(X_{i,t-\tau})),$
where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in range of $0\sim T-1$. Theoretically, infinite time steps are allowed in STBNs, but it is usually finite in practice.
To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as {assumption}[Temporal Assumption] In STBNs, the cause node precedes or parallels the effect node in time. This assumption does not allow the directed cause-effect pairs from now to the past, e.g., $X_{i,t-\tau_{1}}\rightarrow X_{j,t-\tau_{2}},\tau_{1}<\tau_{2}$. And this guarantees the functionally identifiable network structure without Markov equivalence class [6, 21, 88]. Moreover, the causal sufficiency assumption (see Assumption 3.1) is also needed to guarantee the unconfoundedness.
Then, if STBNs are given, the spatial-temporal counterfactuals can be inferred through Algorithm 1, because the global and local STBNs are both DAG. Compared with the naive Bayesian networks, STBNs have the spatio-temporal concept, and the inferring process is naturally forward along with time. This enables to define the momentary interventions for every time steps, and all the interventions would work sequentially, instead of at the same time, like in Algorithm 1. If fully consider this property, the inference algorithm can be faster.

### v3

@@ 4 Differences between POM and SCM
In this section, the fundamental differences between POM and SCM are discussed, focusing on three aspects. (i) Counterfactual falsifiability: POM assumes positivity; therefore, counterfactuals in POM are, in principle, supported by observations and can be empirically falsified. In contrast, counterfactuals in SCM may lie outside the support of the observed data and are thus unfalsifiable. (ii) Ignorability and the back-door criterion: POM requires the ignorability assumption to ensure unconfoundedness and the validity of counterfactual inference. In practice, satisfying ignorability typically requires identifying appropriate covariates. This is a complex, human-guided process. By contrast, SCM achieves unconfoundedness through causal sufficiency and back-door adjustment, which are equivalent to the ignorability assumption. SCM facilitates counterfactual inference through graphical language. (iii) Direct causation and indirect causation: POM struggles to process the indirect causation, and it typically relies on the Baron-Kenny model. In contrast, SCM has the advantage of adjusting for the front-door path (indirect causation) by equivalently applying the back-door criterion twice. We further leverage this advantage to design a forward counterfactual inference algorithm that autonomously infers spatio-temporal counterfactuals. The following are the details.
@@ 4.1 Counterfactual Falsifiability
As shown in Fig. 3, consider the following query, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”. To answer this query, the positivity assumption requires the existence of at least two observational samples that share the same values $X_{1}=0.5$ but differ in $X_{2}$. Specially, one with $X_{2}=1$, and another one with $X_{2}=2$ (see Assumption 2.2). And then, under the consistency assumption (see Assumption 2.2), we can infer counterfactuals by exactly matching the two samples. In contrast to SCM, we can infer the counterfactuals through Pearl’s causal ladder, even if we only have a single sample, as shown in Section 3.1. Thus, how to falsify the counterfactuals?
@@
This is a challenging problem, as counterfactuals are, by definition, unobserved (see “?” in Table 1), and we need at least two samples for comparison. In practice, we cannot return to the past to apply a different treatment, nor can we perfectly replicate the same experiment at a different time. Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [7].
@@
Another explanation is that, without positivity, but with consistency only in POM, counterfactuals can be extrapolated by regressive approaches, and they are falsifiable. For example, as shown in Fig. 5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$. We then query “if had done $T=4$, what would Y=?Y=?”. The counterfactual answer is $Y=2.5$, and the regressive function can falsify the answer because it does not go beyond the observation. Thus, the consistency assumption actually claims that counterfactuals can be falsified by observations. This is also true for SCM, which uses regressive approaches to infer counterfactuals.
@@
Moreover, there may be ambiguity regarding the preset Bayesian network (see Fig. 3). To build an SCM, a Bayesian network must first be known. However, how to discover a probabilistic network from a single sample? This is usually impossible. A reasonable explanation is that the network structure can be inferred from the system’s mechanisms (e.g., physical mechanisms, communication connections) or from graphical knowledge learned from other domains, rather than from probability distribution. However, if it is accepted, the definition of causality could be broader and more ambiguous. This is why probabilistic and causal Bayesian networks are often indistinguishable, as stated in Pearl’s book [7].
@@
Thus, SCM provides a way to think about counterfactuals, emphasizing that the thinking process is counterfactual, not the thinking result, because the result cannot always be falsified without the positivity and consistency assumptions. In contrast, POM emphasizes that the thinking result is counterfactual. This is a fundamental difference in the problem of counterfactual falsifiability.
Figure 4: Diagram of data extrapolation.
Figure 5: Diagram of a back-door path.
@@ 4.2 Ignorability and Back-door Criterion
In the view of statistics, the ignorability assumption (see Assumption 2.2) is equivalent in calculation to the causal sufficiency assumption (see Assumption 3.1) with a back-door criterion. The formal definition of the back-door criterion can be found in these papers and books [17, 18, 4, 5]. Here we present alternative formulation with respect to Table 1, defined as follows:
@@
In Table 1, the covariate $X$ is said to satisfy the back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfies
@@
which is the same in case of $do(T=0)$ [17, 18, 4, 5]. Then, if causal sufficiency holds (i.e., no hidden variables confound $X,Y$ and $T$), the expectation of the potential outcomes in Table 1 can be calculated as
@@
This is the same in POM, as follows:
@@
which is the same in case of $do(T=0)$. Thus, with comparing Eq. (9) and Eq. (10), $\mathbb{E}[Y|do(T=1)]=\mathbb{E}[Y(1)]$, and $\mathbb{E}[Y|do(T=0)]=\mathbb{E}[Y(0)]$, in the two frameworks, but they start from different assumptions. Thus, the ignorability assumption can be decomposed as the causal sufficiency assumption and the back-door criterion, and the causal sufficiency is relatively trivial to satisfy. The ignorability is more likely to be a technical assumption that needs many controlled trials to deconfound, while the back-door path is clear to search in a Bayesian network. A potential, more difficult technical issue is accurately discovering the network structure.
@@ 4.3 Direct and Indirect Causation
Suppose that there is a causal ordered pathway from one treatment variable to another outcome variable. Then, if there is at least one variable between the two endpoints, the pathway is called indirect causation, otherwise, it is called directed causation. Indirect causation also called the mediating effect, has been discussed in depth in Rubin’s papers [74, 75, 76]. Here, the Baron-Kenny model [19, 20] is introduced to present a mediating analysis for a potential outcome using a graph. As shown in Fig. 7, $T\rightarrow Y$ is direct causation, and $T\rightarrow X\rightarrow Y$ is indirect causation.
To analyse indirect causation, the causal-steps approach [19] is proposed by building three linear regression models, as shown in Fig. 7. Suppose that the regressive models are fitted sufficiently, the indirect causation would be detected if satisfy
@@
Another approaches are to test the significance of $H_{0}:b^{\prime}-b=0$ [77, 78, 79]. Or, to test the significance of $H_{0}:ab=0$ [80, 81]. If the correct regressive models are built, counterfactuals can be calculated, just as in SCM in Fig. 3. Moreover, they are all regression-based approaches that are limited for use in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and this process would be seriously time-consuming.
Figure 6: Diagram of mediating analysis [19]. Here, $a,b,c$ are coefficients and $U_{X},U_{Y}$ are biases. $b^{\prime}$ and $U_{total}$ are the coefficient and bias, respectively, for the total effect.
Figure 7: Diagram of front-door adjustment [5]. Here, $T$ is a treatment variable and $Y$ is a target variable. $X_{1}$ and $X_{2}$ are both covariates.
However, the test process can be faster with a Bayesian network in an SCM. Here, we propose an algorithm for forward counterfactual inference, as shown in Algorithm 1. To clearly demonstrate the principle of Algorithm 1, we first introduce the front-door criterion as follows:
@@
The covariate $X$ is said to satisfy the front-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfies
@@
Then, a classic example is provided to show the front-door adjustment using Theorem 3.4.1 in book [5], as shown in Fig. 7 here. If given a treatment variable $T$ and a target variable $Y$, with two covariates $X_{1}$ and $X_{2}$. Then one can adjust the distribution $P(Y|do(T))$ with the front-door criterion as follows:
@@
    If there is one or multiple ordered causal pathways from $T$ to $Y$ and there is at least one mediator on the way, then the causal pathway can be decomposed into two causal pathways through that mediator, e.g., $T\rightarrow X_{1}$ and $X_{1}\rightarrow Y$ in Eq. (11).
@@
Thus, the front-door principle is non-atomic and decomposable. With this more foundational principle in mind, we propose the forward counterfactual algorithm to autonomously apply it, as presented in Algorithm 1. The algorithm can be applied to an arbitrary set of variables $\{X_{1},\dots,X_{n}\}$ that satisfy the causal sufficiency condition. We also provide an example to illustrate the inference process, as shown in Fig. 8.
@@
9      For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to the back-door criterion (see Definition 4.1),
@@
11   end if
12
13 end if
14
15Otherwise, for $q$ pathways starting from $X_{j_{p}}$, select $q$ first mediators $X_{j_{p_{1}}},\dots,X_{j_{p_{q}}}$. Then, a group of source-mediator pairs, e.g., $(X_{j_{p}},X_{j_{p_{q}}})$, can be obtained;
16Factorize $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ as
@@
17Solve each distribution $P(\cdot|do(\cdot),\dots,do(\cdot))$ in Eq. (12) by Algorithm 1 recursively.
@@
Figure 8: Example of the Forward Counterfactual Inference algorithm. Here, $X_{i}$ is the same as $X_{i}^{\prime}$ to distinguish between the two recursions.
In Algorithm 1, the approaches to realize the codes in lines 1 and 3 can be various, e.g., some variants of the depth-first-search algorithm and breadth-first-search algorithm [82]. Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set. Meanwhile, if the computational cost of factorization and adjustment are overlooked, the time complexity of the code in line 9 in Algorithm 1 is $\mathcal{O}(|\mathbf{V}|)$, because we only need to traverse all mediators in each pathway forward, and this process can be performed in parallel. Thus, the total time complexity of Algorithm 1 is $\mathcal{O}(|\mathbf{V}|^{2}+|\mathbf{V}||\mathbf{E}|)$. Thus, if the network structure of the Bayesian network is large and can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
5 Spatio-Temporal Graphical Counterfactuals
Then, for the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns? And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? Would an individual still purchase the product had they not been exposed to the advertisement? In these real-world scenarios, interactive behaviors with lagged causal effects exist widely. POM cannot answer these questions, because it does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 2.2), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [83, 84] and Differences in Differences [85, 86]). This applies to SCM as well, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes. Thus, to answer these counterfactual queries for interactive units and lagged causal effects, we propose a concept of spatio-temporal graphical counterfactuals.
5.1 Spatio-Temporal Bayesian Networks
Before the SCM, various causal graphical models have been proposed to model the spatio-temporal causality. As shown in Fig. 9, a temporal Bayesian network [59, 87] is proposed to model the momentary causal dependency in a first-order Markov process. Further, dynamical Bayesian networks [59] are proposed to represent a combination of multiple temporal Bayesian networks. Full time graph [6] is proposed to summarize temporal Bayesian dependencies with complete timestamps. Full time graph allows the instantaneous causal effects between two units (or variables) in a high-order Markov process, but it only supports strictly stationary causality, that is, causal dependencies that do not change over time. Moreover, the full time graph can be uniquely decomposed into multiple high-order temporal Bayesian networks, named UCN, if the instantaneous causal effects are not allowed [21]. The comparisons of these causal graphs are shown in Table 2.
Figure 9: Diagrams of various causal graphs.
@@ 5.1 Spatio-Temporal Bayesian Networks
Thus, based on the characteristics of these causal graphs, STBNs are defined as a group of DAGs indexed by different timestamps, as shown in Fig. 9. In STBNs, nodes are not equivalent to variables. They represent the state of the variables at some time steps. Under the assumptions of causal Markov (see Assumption 5) and faithfulness (see Assumption 6), the joint distribution of all nodes can be factorized as
$P(X_{1,t},\dots,X_{n,t},X_{1,t-1},\dots,X_{n,t-T+1})=\prod_{\tau=0}^{T-1}\prod_{i=1}^{n}P(X_{i,t-\tau}|\mathbf{Pa}(X_{i,t-\tau})),$ (13)
where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in the range of $0\sim T-1$. Theoretically, STBNs allow an infinite number of time steps, but in practice, they are typically finite.
To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as {assumption}[Temporal Assumption] In STBNs, the cause node precedes or occurs concurrently with the effect node. This assumption does not allow the directed cause-effect pairs from now to the past, e.g., $X_{i,t-\tau_{1}}\rightarrow X_{j,t-\tau_{2}},\tau_{1}<\tau_{2}$. And this guarantees a functionally identifiable network structure without Markov equivalence classes [6, 21, 88]. Moreover, the causal sufficiency assumption (see Assumption 3.1) is also needed to guarantee the unconfoundedness.
Then, if STBNs are given, the spatial-temporal counterfactuals can be inferred through Algorithm 1, because the global and local STBNs are both DAGs. Compared with naive Bayesian networks, STBNs incorporate spatio-temporal concepts, and the inference process naturally proceeds forward in time. This enables momentary interventions at every time step, and all interventions work sequentially rather than simultaneously, as in Algorithm 1. If this property is considered thoroughly, the inference algorithm can be faster.